# Functional Central Limit Theorem for the simultaneous subgraph count of dynamic Erdős-Rényi random graphs

### v1

@@ 4. Main result
$\bar{X}_{N,i}(t)=\frac{X^{\star}_{N,i}(t)}{\bigl{(}p_{\star}(t)\bigr{)}^{\mathcal{E}({H}_{i})}}$
Due to Proposition 5.3, to obtain the assertion of Lemma 5.4, it is enough to show that the vector $\bigl{(}\bar{\boldsymbol{X}}_{N}(t_{1}),\ldots,\bar{\boldsymbol{X}}_{N}(t_{n})\bigr{)}$ converges in distributions to a centered Gaussian random vector with a covariance block matrix $\bar{\boldsymbol{\Sigma}}$ with the entries
@@
for any $k,l\in\langle n\rangle$ and any $i,j\in\langle m\rangle$. To do so we use the Cramér-Wold device: it is enough to show that for any vectors $\boldsymbol{q}_{1},\ldots\boldsymbol{q}_{n}\in\mathbb{R}^{m}$ random variable
@@
$\lim_{N\to\infty}\mathbb{E}\left\{\xi_{{N}}^{z}\right\}=\begin{cases}\sigma^{z}(z-1)!!,\qquad\text{if $z$ is even},\\ 0,\qquad\text{if $z$ is odd}\end{cases}$ (7.3)
@@
$\xi_{i,{N}}(g)=\sum_{k=1}^{n}\frac{q_{k,i}\bigl{(}\mathbb{I}\{g{\,\sqsubset\,}G_{N}(t_{k})\}-\mathbb{P}\left\{g{\,\sqsubset\,}G_{N}(t_{k})\right\}\bigr{)}}{\bigl{(}p_{\star}(t_{k})\bigr{)}^{\mathcal{E}({H}_{i})}}.$ (7.5)
@@
[…]
@@
We now move to the next step of our proof: identifying the terms in the inner sum of (7.10) that determine the asymptotics of this sum (i.e., the remaining terms are asymptotically negligible). To do so, classify all the collections $\boldsymbol{g}$ that appear in (7.10) by the number of vertices in their union, i.e., the value $\mathcal{V}(\boldsymbol{g}):=\mathcal{V}\left(\bigcup_{i=1}^{z}g_{i}\right)$. In this case we can first choose a complete graph $K_{\mathcal{V}(\boldsymbol{g})}{\,\sqsubset\,}K_{N}$ where the graph $\bigcup_{i=1}^{z}g_{i}$ is located (where it is noted that the number of such choices is dependent on $N$), after which we choose $g_{i}{\,\sqsubset\,}K_{\mathcal{V}(\boldsymbol{g})}$ (where the number of such choices no longer depends neither on $N$, nor on the particular choice of the graph $K_{\mathcal{V}(\boldsymbol{g})}$). Let
$V_{\max}=\max\limits_{\boldsymbol{g}\in{\tt G}^{\star}_{\boldsymbol{f},N}}\mathcal{V}(\boldsymbol{g}).$
@@
$=\sum_{{v=2}}^{V_{\max}}\sum_{K_{v}{\,\sqsubset\,}K_{N}}\sum_{\begin{subarray}{c}\boldsymbol{g}\in{\tt G}^{\star}_{\boldsymbol{f},v}\\ \mathcal{V}\left(\boldsymbol{g}\right)=v\end{subarray}}\mathbb{E}\left\{\prod_{i=1}^{z}\frac{\mathcal{F}_{{N}}({H}_{f_{i}})\xi_{f_{i},{N}}(g_{i})}{\sqrt{\mathcal{F}^{\rm opt}_{N}({H}_{f_{i}})}}\right\}$
@@
Continuation of proof of Lemma 5.4. Recall that from (7.9) we know that, for any $\boldsymbol{g}\in{{\tt G}^{\star}_{\boldsymbol{f},v}}$ each connected component of $\mathcal{G}(\boldsymbol{g})$ has at least two vertices. Thus, if the graph $\mathcal{G}(\boldsymbol{g})$ contains at least one connected component with at least three vertices, then due to Lemma 7.1, in combination with (7.10) and (7.11), we obtain that
@@
In addition, for even $z$, we restrict our attention to the ${\color[rgb]{0,0,1}\boldsymbol{g}}\in{\tt G}_{\boldsymbol{f},v}$ for which every connected component of $\mathcal{G}(\boldsymbol{g})$ consists of exactly two vertices, meaning that it forms a ‘perfect matching’ (again, because $\mathcal{G}(\boldsymbol{g})$ has no isolated vertices).
@@
$\mathbb{E}\left\{N^{\mathcal{V}({\color[rgb]{0,0,1}g_{i}}\cup g_{j})}\frac{\mathcal{F}_{{N}}({H}_{f_{i}})\xi_{f_{i},{N}}(g_{i})}{\sqrt{\mathcal{F}^{\rm opt}_{N}({H}_{f_{i}})}}\frac{\mathcal{F}_{{N}}({H}_{f_{j}})\xi_{f_{j},{N}}(g_{j})}{\sqrt{\mathcal{F}^{\rm opt}_{N}({H}_{f_{j}})}}\right\}$ $=\frac{N^{\mathcal{V}(g_{i}\cup g_{j})}\mathcal{F}_{{N}}({H}_{f_{i}})\mathcal{F}_{{N}}({H}_{f_{j}})}{\sqrt{\mathcal{F}^{\rm opt}_{N}({H}_{f_{i}})}\sqrt{\mathcal{F}^{\rm opt}_{N}({H}_{f_{j}})}}$
×$\times\sum_{k,l=1}^{n}\frac{q_{k,i}q_{l,j}\operatorname{Cov}\bigl{(}\mathbb{I}\{g_{i}{\,\sqsubset\,}G_{N}(t_{k})\},\mathbb{I}\{g_{j}{\,\sqsubset\,}G_{N}(t_{l})\}\bigr{)}}{\bigl{(}p_{\star}(t_{k})\bigr{)}^{\mathcal{E}({H}_{i})}\bigl{(}p_{\star}(t_{l})\bigr{)}^{\mathcal{E}({H}_{j})}}.$ (7.14)
@@
×$\qquad\times\bigl{(}p_{\star}({t_{l}})\bigr{)}^{\mathcal{E}({H}_{j})-\mathcal{E}(g_{i}\cap g_{j})}\bigl{(}\kappa_{\star}({t_{k},t_{l}})\bigr{)}^{\mathcal{E}(g_{i}\cap g_{j})}.$
@@
×$\qquad\times\varrho_{N}^{-\mathcal{E}(g_{i})}\varrho_{N}^{-\mathcal{E}(g_{j})}\varrho_{N}^{\mathcal{E}({H}_{i})+\mathcal{E}({H}_{j})-\mathcal{E}(g_{1}\cap g_{2})}$
@@
$={\bigl{(}1+o(1)\bigr{)}}\sqrt{\frac{\mathcal{F}_{N}(g_{i}\cap g_{j})}{\mathcal{F}^{\rm{opt}}_{N}({H}_{f_{i}})}}\sqrt{\frac{\mathcal{F}_{N}(g_{i}\cap g_{j})}{\mathcal{F}^{\rm{opt}}_{N}({H}_{f_{j}})}}\times\sum_{k,l=1}^{n}\frac{q_{k,i}q_{l,j}\bigl{(}\kappa_{\star}({t_{k},t_{l}})\bigr{)}^{\mathcal{E}(g_{i}\cap g_{j})}}{\bigl{(}p_{\star}({t_{k}})\,p_{\star}({t_{l}})\bigr{)}^{\mathcal{E}(g_{i}\cap g_{j})}}.$ (7.15)
Recall that from (3.10), as $g_{i}\cap g_{j}{\,\sqsubset\,}g_{i}\sim{H}_{f_{i}}$, we know that
$\lim_{N\to\infty}\sqrt{\frac{\mathcal{F}_{N}(g_{i}\cap g_{j})}{\mathcal{F}^{\rm{opt}}_{N}({H}_{f_{i}})}}=\mathbb{I}\{g_{i}\cap g_{j}\in{\tt OCS}({H}_{f_{i}})\}.$
@@
$\lim_{N\to\infty}\sqrt{\frac{\mathcal{F}_{N}(g_{i}\cap g_{j})}{\mathcal{F}^{\rm{opt}}_{N}({H}_{f_{j}})}}=\mathbb{I}\{g_{i}\cap g_{j}\in{\tt OCS}({H}_{f_{j}})\}.$
@@
$=\mathbb{I}\{{H}_{f_{i}}\cong{H}_{f_{j}}\}\mathbb{I}\{g_{i}\cap g_{j}\in{\tt OCS}({H}_{f_{i}},{H}_{f_{j}})\}\times\sum_{k,l=1}^{n}\frac{q_{k,i}q_{l,j}\bigl{(}\kappa_{\star}({t_{k},t_{l}})\bigr{)}^{\mathcal{E}(g_{i}\cap g_{j})}}{\bigl{(}p_{\star}({t_{k}})p_{\star}({t_{l}})\bigr{)}^{\mathcal{E}(g_{i}\cap g_{j})}}.$ (7.16)
@@
${\mathbb{J}}({\boldsymbol{f}},{\boldsymbol{g}},H)$ $=\prod_{(i,j)\in{\tt E}({H})}\mathbb{I}\{{H}_{f_{i}}\cong{H}_{f_{j}}\}\Bigg{(}\mathbb{I}\{g_{i}\cap g_{j}\in{\tt OCS}({H}_{f_{i}},{H}_{f_{j}})\}\sum_{k,l=1}^{n}\frac{q_{k,i}q_{l,j}\bigl{(}\kappa_{\star}(t_{k},t_{l})\bigr{)}^{\mathcal{E}(g_{i}\cap g_{j})}}{\bigl{(}p_{\star}(t_{k})\,p_{\star}(t_{l})\bigr{)}^{\mathcal{E}(g_{i}\cap g_{j})}}\Bigg{)}.$
@@
Note that each $\boldsymbol{g}\in{\tt G}_{\boldsymbol{f}}^{\star}$ from the third sum in (7.17) corresponds uniquely to a specific $v\in\langle V_{\max}\rangle$ in the second sum (where $v=\mathcal{V}(\boldsymbol{g})$). Consequently, the following two sums are equivalent:
$\sum_{v=0}^{V_{\max}}\sum_{\begin{subarray}{c}\boldsymbol{g}\in{\tt G}^{\star}_{\boldsymbol{f}}\\ \mathcal{V}\left(\boldsymbol{g}\right)=v\\ \mathcal{G}(\boldsymbol{g})={H}\end{subarray}}\equiv\sum_{\begin{subarray}{c}\boldsymbol{g}\in{\tt G}^{\star}_{\boldsymbol{f}}\\ \mathcal{G}(\boldsymbol{g})={H}\end{subarray}}.$
@@
Denote the edges of graph ${H}\in{\tt M}_{z}$ as follows: ${\tt E}({H})=\{(i_{1},j_{1}),\ldots,(i_{z/2},j_{z/2})\}$. Note that the right-hand side of (7.18) does not depend on the full structure of vector $\boldsymbol{g}$, but only on the structures of graphs $g_{i_{b}}\cap g_{j_{b}}$, for $b\in\langle z/2\rangle$. Hence we decompose this sum by considering the various types of ‘structures’ of the graphs $g_{i_{b}}\cap g_{j_{b}}$, noting that due to the indicators appearing in the right-hand side of (7.18) we can restrict ourselves to ${g_{i_{b}}\cap g_{j_{b}}}\in{\tt OCS}(H_{f_{i_{b}}},H_{f_{j_{b}}})$ only, so as to obtain
@@
${\tt G}^{\star\star}_{\boldsymbol{f}}({H},\boldsymbol{g}^{\cap}):=\left\{\boldsymbol{g}\in{\tt G}_{\boldsymbol{f}}^{\star}\colon\mathcal{G}(\boldsymbol{g})={H},\,\forall b\in\langle z/2\rangle:g_{i_{b}}\cap g_{j_{b}}\sim g^{\cap}_{b}\right\},$
@@
×$\qquad\qquad\qquad\times\prod_{b=1}^{z/2}\sum_{k,l=1}^{n}\frac{q_{k,i}q_{l,j}\bigl{(}\kappa_{\star}({t_{k},t_{l}})\bigr{)}^{\mathcal{E}(g_{b}^{\cap})}}{\bigl{(}p_{\star}({t_{k}})\,p_{\star}({t_{l}})\bigr{)}^{\mathcal{E}(g_{b}^{\cap})}}$
@@
×$\qquad\qquad\qquad\times\prod_{b=1}^{z/2}\sum_{k,l=1}^{n}\frac{q_{k,i}q_{l,j}\bigl{(}\kappa_{\star}({t_{k},t_{l}})\bigr{)}^{\mathcal{E}(g_{b}^{\cap})}}{\bigl{(}p_{\star}({t_{k}})\,p_{\star}({t_{l}})\bigr{)}^{\mathcal{E}(g_{b}^{\cap})}}.$
@@
Continuation of proof of Lemma 5.4. Combining (7.20) with the result obtained in Lemma 7.2, we can now derive the asymptotics of the inner sum of (7.10):
${\lim_{N\to\infty}}\sum_{\boldsymbol{g}\in{\tt G}^{\star}_{\boldsymbol{f},N}}\mathbb{E}\left\{\prod_{i=1}^{z}\frac{\mathcal{F}_{{N}}({H}_{f_{i}})\xi_{f_{i},{N}}(g_{i})}{\sqrt{\mathcal{F}^{\rm opt}_{N}({H}_{f_{i}})}}\right\}$ $=\sum_{{H}\in{\tt M}_{z}}\left(\prod_{b=1}^{z/2}\mathbb{I}\{{H}_{f_{i_{b}}}\cong{H}_{f_{j_{b}}})\}\right)$
×$\qquad\times\sum_{\boldsymbol{g}^{\cap}\in{\tt OCS}_{\boldsymbol{f}}({H})}\prod_{b=1}^{z/2}\mathcal{C}({H}_{f_{i_{b}}},{H}_{f_{j_{b}}},g^{\cap}_{b})\sum_{k,l=1}^{n}\frac{q_{k,i}q_{l,j}\bigl{(}\kappa_{\star}(k,l)\bigr{)}^{\mathcal{E}(g_{b}^{\cap})}}{\bigl{(}p_{\star}(k)p_{\star}(l)\bigr{)}^{\mathcal{E}(g_{b}^{\cap})}}$
$=\sum_{{H}\in{\tt M}_{z}}\left(\prod_{b=1}^{z/2}\mathbb{I}\{{H}_{f_{i_{b}}}\cong{H}_{f_{j_{b}}})\}\right)$
×$\qquad\times\prod_{b=1}^{z/2}\left(\sum_{k,l=1}^{n}q_{k,i}q_{l.j}\sum_{g^{\cap}\in{\tt OCS}({H}_{f_{i_{b}}},{H}_{f_{j_{b}}})}\frac{\mathcal{C}({H}_{f_{i_{b}}},{H}_{f_{j_{b}}},g^{\cap})\bigl{(}\kappa_{\star}(k,l)\bigr{)}^{\mathcal{E}(g^{\cap})}}{\bigl{(}p_{\star}(k)p_{\star}(l)\bigr{)}^{\mathcal{E}(g^{\cap})}}\right)$
@@
$=\sum_{{H}\in{\tt M}_{z}}\sum_{f_{1},\ldots,f_{z}\in\langle m\rangle}\prod_{b=1}^{z/2}\left(\sum_{k,l=1}^{n}q_{k,i}q_{l.j}\left(\Sigma_{k,l}\right)_{f_{i_{b}},f_{j_{b}}}\right)=\sum_{{H}\in{\tt M}_{z}}\left(\sum_{i,j=1}^{m}\sum_{k,l=1}^{n}q_{k,i}q_{l,j}\left(\Sigma_{k,l}\right)_{i,j}\right)^{z/2}$

### v2

@@ 7 Proof of Lemma 5
$\bar{X}_{N,i}(t)=\frac{X^{\star}_{N,i}(t)}{\bigl(p_{\star}(t)\bigr)^{\mathcal{E}({H}_{i})}}{.}$
Due to Proposition 4, to obtain the assertion of Lemma 5, it is enough to show that the vector $\bigl(\bar{\boldsymbol{X}}_{N}(t_{1}),\ldots,\bar{\boldsymbol{X}}_{N}(t_{n})\bigr)$ converges in distributions to a centered Gaussian random vector with a covariance block matrix $\bar{\boldsymbol{\Sigma}}$ with the entries
@@
for any $k,l\in\langle n\rangle$ and any $i,j\in\langle m\rangle$. To do so we use the Cramér–Wold device: it is enough to show that, for any vectors $\boldsymbol{q}_{1},\ldots\boldsymbol{q}_{n}\in\mathbb{R}^{m}$, the random variable
@@
${\lim_{N\to\infty}\mathbb{E}\left\{\xi_{{N}}^{z}\right\}=\begin{cases}\sigma^{z}(z-1)!!,\qquad&\text{if $z$ is even},\\ 0,\qquad&\text{if $z$ is odd,}\end{cases}}$ (7.3)
@@
$\xi_{i,{N}}(g)=\sum_{k=1}^{n}\frac{q_{k,i}\bigl(\mathbb{I}\{g{\,{\subset}\,}G_{N}(t_{k})\}-\mathbb{P}\left\{g{\,{\subset}\,}G_{N}(t_{k})\right\}\bigr)}{\bigl(p_{\star}(t_{k})\bigr)^{\mathcal{E}({H}_{i})}}.$ (7.5)
@@
Then, the last sum in (7.7) is simply a sum over $\boldsymbol{g}=(g_{1},\ldots,g_{z})^{\top}\in{\tt G}_{\boldsymbol{f},N}$}. The primary objective of this proof is to demonstrate that (7.7) satisfies the asymptotics given by (7.3). The first step involves identifying the terms in the inner sum of (7.7) that are non-zero. Next, we isolate the terms within that sum which make the dominant contribution to the asymptotics. Finally, we compute the asymptotics of the identified dominant terms. We start the our first step by investigating for which collections $\boldsymbol{g}$ the expectation at the end of (7.7) is different from zero. Using (3.2) we observe that if $\mathcal{E}(g^{\star}\cap g^{\star\star})=0$ for some subgraphs $g^{\star},g^{\star\star}{\,{\subset}\,}K_{N}$, then, for any $i,i^{\star}\in\langle m\rangle$ the random variables $\xi_{i,{N}}(g^{\star})$ and $\xi_{i^{\star},{N}}(g^{\star\star})$ are independent. Hence, if for some $j\in\langle z\rangle$ and all $i\in\langle k\rangle$, $i\not=j$ we have that $\mathcal{E}(g_{i}\cap g_{j})=0$, then
@@
We now move to the next step of our proof: identifying the terms in the inner sum of (7.10) that determine the asymptotics of this sum (i.e., the remaining terms are asymptotically negligible). To do so, classify all the collections $\boldsymbol{g}$ that appear in (7.10) by the number of vertices in their union, i.e., the value $\mathcal{V}(\boldsymbol{g}):=\mathcal{V}\left(\bigcup_{i=1}^{z}g_{i}\right)$. In this case we can first choose a complete graph $K_{\mathcal{V}(\boldsymbol{g})}{\,{\subset}\,}K_{N}$ where the graph $\bigcup_{i=1}^{z}g_{i}$ is located (where it is noted that the number of such choices is dependent on $N$), after which we choose $g_{i}{\,{\subset}\,}K_{\mathcal{V}(\boldsymbol{g})}$ (where the number of such choices no longer depends neither on $N$, nor on the particular choice of the graph $K_{\mathcal{V}(\boldsymbol{g})}$). Let
$V_{\max}=\max\limits_{\boldsymbol{g}\in{\tt G}^{\star}_{\boldsymbol{f},N}}\mathcal{V}(\boldsymbol{g}).$ (7.11)
@@
$=\sum_{{v=2}}^{V_{\max}}\sum_{K_{v}{\,{\subset}\,}K_{N}}\sum_{\begin{subarray}{c}\boldsymbol{g}\in{\tt G}^{\star}_{\boldsymbol{f},v}\\ \mathcal{V}\left(\boldsymbol{g}\right)=v\end{subarray}}\mathbb{E}\left\{\prod_{i=1}^{z}\frac{\mathcal{F}_{{N}}({H}_{f_{i}})\xi_{f_{i},{N}}(g_{i})}{\sqrt{\mathcal{F}^{\rm opt}_{N}({H}_{f_{i}})}}\right\}$
@@
Continuation of proof of Lemma 5. Recall that from (7.9) we know that, for any $\boldsymbol{g}\in{{\tt G}^{\star}_{\boldsymbol{f},v}}$ each connected component of $\mathcal{G}(\boldsymbol{g})$ has at least two vertices. Thus, if the graph $\mathcal{G}(\boldsymbol{g})$ contains at least one connected component with at least three vertices, then due to Lemma 7, in combination with (7.10) and (7.12), we obtain that
@@
In addition, for even $z$, we restrict our attention to the ${\boldsymbol{g}}\in{\tt G}_{\boldsymbol{f},v}$ for which every connected component of $\mathcal{G}(\boldsymbol{g})$ consists of exactly two vertices, meaning that it forms a ‘perfect matching’ (again, because $\mathcal{G}(\boldsymbol{g})$ has no isolated vertices).
@@
$\mathbb{E}\left\{N^{\mathcal{V}({g_{i}}\cup g_{j})}\frac{\mathcal{F}_{{N}}({H}_{f_{i}})\xi_{f_{i},{N}}(g_{i})}{\sqrt{\mathcal{F}^{\rm opt}_{N}({H}_{f_{i}})}}\frac{\mathcal{F}_{{N}}({H}_{f_{j}})\xi_{f_{j},{N}}(g_{j})}{\sqrt{\mathcal{F}^{\rm opt}_{N}({H}_{f_{j}})}}\right\}$ $=\frac{N^{\mathcal{V}(g_{i}\cup g_{j})}\mathcal{F}_{{N}}({H}_{f_{i}})\mathcal{F}_{{N}}({H}_{f_{j}})}{\sqrt{\mathcal{F}^{\rm opt}_{N}({H}_{f_{i}})}\sqrt{\mathcal{F}^{\rm opt}_{N}({H}_{f_{j}})}}$
$\hskip-28.45274pt\times\sum_{k,l=1}^{n}\frac{q_{k,i}q_{l,j}\operatorname{Cov}\bigl(\mathbb{I}\{g_{i}{\,{\subset}\,}G_{N}(t_{k})\},\mathbb{I}\{g_{j}{\,{\subset}\,}G_{N}(t_{l})\}\bigr)}{\bigl(p_{\star}(t_{k})\bigr)^{\mathcal{E}({H}_{i})}\bigl(p_{\star}(t_{l})\bigr)^{\mathcal{E}({H}_{j})}}.$ (7.15)
@@
$\qquad\times\bigl(p_{\star}({t_{l}})\bigr)^{\mathcal{E}({H}_{j})-\mathcal{E}(g_{i}\cap g_{j})}\bigl(\kappa_{\star}({t_{k},t_{l}})\bigr)^{\mathcal{E}(g_{i}\cap g_{j})}.$
@@
$\qquad\times\varrho_{N}^{-\mathcal{E}(g_{i})}\varrho_{N}^{-\mathcal{E}(g_{j})}\varrho_{N}^{\mathcal{E}({H}_{i})+\mathcal{E}({H}_{j})-\mathcal{E}(g_{1}\cap g_{2})}$
@@
$\hskip-186.36563pt={\bigl(1+o(1)\bigr)}\sqrt{\frac{\mathcal{F}_{N}(g_{i}\cap g_{j})}{\mathcal{F}^{\rm{opt}}_{N}({H}_{f_{i}})}}\sqrt{\frac{\mathcal{F}_{N}(g_{i}\cap g_{j})}{\mathcal{F}^{\rm{opt}}_{N}({H}_{f_{j}})}}\times\sum_{k,l=1}^{n}\frac{q_{k,i}q_{l,j}\bigl(\kappa_{\star}({t_{k},t_{l}})\bigr)^{\mathcal{E}(g_{i}\cap g_{j})}}{\bigl(p_{\star}({t_{k}})\,p_{\star}({t_{l}})\bigr)^{\mathcal{E}(g_{i}\cap g_{j})}}.$ (7.16)
Recall that from (3.9), as $g_{i}\cap g_{j}{\,{\subset}\,}g_{i}\sim{H}_{f_{i}}$, we know that
$\lim_{N\to\infty}\sqrt{\frac{\mathcal{F}_{N}(g_{i}\cap g_{j})}{\mathcal{F}^{\rm{opt}}_{N}({H}_{f_{i}})}}=\mathbb{I}\{{\pi(g_{i}\cap g_{j})\in\,}{\tt OCS}({H}_{f_{i}})\}.$
@@
$\lim_{N\to\infty}\sqrt{\frac{\mathcal{F}_{N}(g_{i}\cap g_{j})}{\mathcal{F}^{\rm{opt}}_{N}({H}_{f_{j}})}}=\mathbb{I}\{{\pi(g_{i}\cap g_{j})\in\,}{\tt OCS}({H}_{f_{j}})\}.$
@@
$\hskip-227.62204pt=\mathbb{I}\{{H}_{f_{i}}\cong{H}_{f_{j}}\}\mathbb{I}\{{\pi(g_{i}\cap g_{j})\in\,}{\tt OCS}({H}_{f_{i}},{H}_{f_{j}})\}\times\sum_{k,l=1}^{n}\frac{q_{k,i}q_{l,j}\bigl(\kappa_{\star}({t_{k},t_{l}})\bigr)^{\mathcal{E}(g_{i}\cap g_{j})}}{\bigl(p_{\star}({t_{k}})p_{\star}({t_{l}})\bigr)^{\mathcal{E}(g_{i}\cap g_{j})}}.$ (7.17)
@@
${\mathbb{J}}({\boldsymbol{f}},{\boldsymbol{g}},H)$ $=\prod_{(i,j)\in{\tt E}({H})}\mathbb{I}\{{H}_{f_{i}}\cong{H}_{f_{j}}\}\Bigg(\mathbb{I}\{{\pi(g_{i}\cap g_{j})\in\,}{\tt OCS}({H}_{f_{i}},{H}_{f_{j}})\}\sum_{k,l=1}^{n}\frac{q_{k,i}q_{l,j}\bigl(\kappa_{\star}(t_{k},t_{l})\bigr)^{\mathcal{E}(g_{i}\cap g_{j})}}{\bigl(p_{\star}(t_{k})\,p_{\star}(t_{l})\bigr)^{\mathcal{E}(g_{i}\cap g_{j})}}\Bigg).$
@@
Note that each $\boldsymbol{g}\in{\tt G}_{\boldsymbol{f}}^{\star}$ from the third sum in (7.18) corresponds uniquely to a specific $v\in\langle V_{\max}\rangle$ in the second sum (where $v=\mathcal{V}(\boldsymbol{g})$). Consequently, recalling the definition of $V_{\max}$ given in (7.11),
$\bigcup_{v=1}^{V_{\max}}\left\{\boldsymbol{g}\in{\tt G}^{\star}_{\boldsymbol{f}}\colon\mathcal{V}(\boldsymbol{g})=v,\,\mathcal{G}(\boldsymbol{g})=H\right\}$ $=\left\{\boldsymbol{g}\in{\tt G}_{\boldsymbol{f}}^{\star}\colon\mathcal{G}(\boldsymbol{g})=H,\,\mathcal{V}(\boldsymbol{g})\in\langle V_{\max}\rangle\right\}$
$=\left\{\boldsymbol{g}\in{\tt G}_{\boldsymbol{f}}^{\star}\colon\mathcal{G}(\boldsymbol{g})=H\right\},$
and for any $v_{1},v_{2}\in\langle V_{\max}\rangle$, $v_{1}\not=v_{2}$
$\left\{\boldsymbol{g}\in{\tt G}^{\star}_{\boldsymbol{f}}\colon\mathcal{V}(\boldsymbol{g})=v_{1},\,\mathcal{G}(\boldsymbol{g})=H\right\}\cap\left\{\boldsymbol{g}\in{\tt G}^{\star}_{\boldsymbol{f}}\colon\mathcal{V}(\boldsymbol{g})=v_{2},\,\mathcal{G}(\boldsymbol{g})=H\right\}=\varnothing,$
implying that
$\sum_{v=0}^{V_{\max}}\sum_{\begin{subarray}{c}\boldsymbol{g}\in{\tt G}^{\star}_{\boldsymbol{f}}\\ \mathcal{V}\left(\boldsymbol{g}\right)=v\\ \mathcal{G}(\boldsymbol{g})={H}\end{subarray}}\frac{1}{\mathcal{V}(\boldsymbol{g})!}{\mathbb{J}}({\boldsymbol{f}},{\boldsymbol{g}},H)=\sum_{\begin{subarray}{c}\boldsymbol{g}\in{\tt G}^{\star}_{\boldsymbol{f}}\\ \mathcal{G}(\boldsymbol{g})={H}\end{subarray}}\frac{1}{\mathcal{V}(\boldsymbol{g})!}{\mathbb{J}}({\boldsymbol{f}},{\boldsymbol{g}},H).$
@@
Denote the edges of graph ${H}\in{\tt M}_{z}$ as follows: ${\tt E}({H})=\{(i_{1},j_{1}),\ldots,(i_{z/2},j_{z/2})\}$. Note that the right-hand side of (7.19) does not depend on the full structure of vector $\boldsymbol{g}$, but only on the patterns of graphs $g_{i_{b}}\cap g_{j_{b}}$, for $b\in\langle z/2\rangle$. Hence we decompose this sum by considering the various patterns $g_{b}^{\cap}$ of the graphs $g_{i_{b}}\cap g_{j_{b}}$, noting that due to the indicators appearing in the right-hand side of (7.19) we can restrict ourselves to ${g_{b}^{\cap}\in}{\tt OCS}(H_{f_{i_{b}}},H_{f_{j_{b}}})$ only, so as to obtain
@@
${\tt G}^{\star\star}_{\boldsymbol{f}}({H},\boldsymbol{g}^{\cap}):=\left\{\boldsymbol{g}\in{\tt G}_{\boldsymbol{f}}^{\star}\colon\mathcal{G}(\boldsymbol{g})={H},\,\forall b\in\langle z/2\rangle:{\pi(g_{i_{b}}\cap g_{j_{b}})=}g^{\cap}_{b}\right\},$
@@
$\qquad\qquad\qquad\times\prod_{b=1}^{z/2}\sum_{k,l=1}^{n}\frac{q_{k,i}q_{l,j}\bigl(\kappa_{\star}({t_{k},t_{l}})\bigr)^{\mathcal{E}(g_{b}^{\cap})}}{\bigl(p_{\star}({t_{k}})\,p_{\star}({t_{l}})\bigr)^{\mathcal{E}(g_{b}^{\cap})}}$
@@
$\qquad\qquad\qquad\times\prod_{b=1}^{z/2}\sum_{k,l=1}^{n}\frac{q_{k,i}q_{l,j}\bigl(\kappa_{\star}({t_{k},t_{l}})\bigr)^{\mathcal{E}(g_{b}^{\cap})}}{\bigl(p_{\star}({t_{k}})\,p_{\star}({t_{l}})\bigr)^{\mathcal{E}(g_{b}^{\cap})}}.$
@@
Continuation of proof of Lemma 5. Combining (7.21) with the result obtained in Lemma 8, we can now derive the asymptotics of the inner sum of (7.10):
$=\sum_{{H}\in{\tt M}_{z}}\left(\prod_{b=1}^{z/2}\mathbb{I}\{{H}_{f_{i_{b}}}\cong{H}_{f_{j_{b}}})\}\right)\times\sum_{\boldsymbol{g}^{\cap}\in{\tt OCS}_{\boldsymbol{f}}({H})}\prod_{b=1}^{z/2}\mathcal{C}({H}_{f_{i_{b}}},{H}_{f_{j_{b}}},g^{\cap}_{b})\sum_{k,l=1}^{n}\frac{q_{k,i}q_{l,j}\bigl(\kappa_{\star}(k,l)\bigr)^{\mathcal{E}(g_{b}^{\cap})}}{\bigl(p_{\star}(k)p_{\star}(l)\bigr)^{\mathcal{E}(g_{b}^{\cap})}}$
$=\sum_{{H}\in{\tt M}_{z}}\left(\prod_{b=1}^{z/2}\mathbb{I}\{{H}_{f_{i_{b}}}\cong{H}_{f_{j_{b}}})\}\right)$ (7.22)
$\qquad\times\prod_{b=1}^{z/2}\left(\sum_{k,l=1}^{n}q_{k,i}q_{l.j}\sum_{g^{\cap}\in{\tt OCS}({H}_{f_{i_{b}}},{H}_{f_{j_{b}}})}\frac{\mathcal{C}({H}_{f_{i_{b}}},{H}_{f_{j_{b}}},g^{\cap})\bigl(\kappa_{\star}(k,l)\bigr)^{\mathcal{E}(g^{\cap})}}{\bigl(p_{\star}(k)p_{\star}(l)\bigr)^{\mathcal{E}(g^{\cap})}}\right)$
@@
$=\sum_{{H}\in{\tt M}_{z}}\sum_{f_{1},\ldots,f_{z}\in\langle m\rangle}\prod_{b=1}^{z/2}\left(\sum_{k,l=1}^{n}q_{k,i}q_{l.j}\left(\Sigma_{k,l}\right)_{f_{i_{b}},f_{j_{b}}}\right)$
$=\sum_{{H}\in{\tt M}_{z}}\left(\sum_{i,j=1}^{m}\sum_{k,l=1}^{n}q_{k,i}q_{l,j}\left(\Sigma_{k,l}\right)_{i,j}\right)^{z/2}$